OpenStudy (anonymous):

A jar contains 7 blue cubes, 4 blue spheres, 5 green cubes, and 6 green spheres. If you select an object at random, what is the probability that the object is green or a cube?

OpenStudy (anonymous):

Add all the stuff up: 7 + 4 + 5 + 6 = 11 + 11 = 22 probability of a green object or a cube, so add the green objects and the total cubes 11 green objects + 12 cubes. = 23 23 / 22

OpenStudy (anonymous):

it is an OR operation. 1) A = event that the object is green P(A) = (5+6)/(7+4+5+6) P(A) = 11/22 P(A) = 1/2 2) B = event that the object is a cube P(B) = (7+5)/(7+4+5+6) P(B) = 12/22 P(B) = 6/11 3) A or B P(A U B) = P(A) + P(B) - P(A intersection B)

OpenStudy (anonymous):

@callmemaybe16 probability shoul;d never be more than "1"

OpenStudy (anonymous):

4) we now need that intersection probability i.e., probability of both happening. P(A and B) = 5/22 so, going back to (3) P(A or B) = (23/22) - (5/22) = 18/22